On December 31, 2017, Roper Company had current assets (cash) of $15,000 and current liabilities (accounts payable) of $8,000, resulting in a current ratio of 1.88. The company needs to increase its current ratio to 2.75 by December 31, 2018. Calculate the amount of accounts payable that needs to be paid in order to boost the current ratio to 2.75.
Answer #1

Current ratio formula is

current assets/current liabilities

So present current ratio is 1.88

Current assets /current liabilities =15000/8000=1.88

In order to boost current ratio to 2.75 now without increasing the assets then current liabilities need to paid off to some EXTENT.as result current asset(cash gets reduced to the extent of payment made to payables)

LET X AMOUNT NEED TO BE PAID OFF to payables

THEN

EQUATION WILL BE current assets/current liabilities

=(15000-X) /(8000-X) =2.75

=(15000-X) =2.75(8000-X)

=15000-X=22000-2.75X

=2.75X-X=22000-15000

=X=7000/1.75

=X =40000

Hence the amount to be paid to payable is 4000

Now current ratio will be 15000-4000/8000-4000=2.75 Times.